how do you graph y=3x-2
Answer by TimothyLamb(4379)

You need two points on the line in order to graph it.
Plug-in two different x values in the equation to get their y values, like so:
---
Point p1, use x=0:
y = 3(0) - 2
y = -2
p1 = (0, -2)
---
Point p2, use x=1:
y = 3(1) - 2
y = 1
p2 = (1, 1)
---
Now draw a line between points p1 and p2 to graph the equation.
The line extends infinitely beyond both points.
---
To check the answer, calculate the slope from the two points, like so:
Calculate slope: m = (y2 - y1)/(x2 - x1)
m = (1 - -2) / (1 - 0)
m = 3 / 1
m = 3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php